# Efficient Beam Selection for ISAC in Cell-Free Massive MIMO via Digital Twin-Assisted Deep Reinforcement Learning

### v1

@@ IV Digital Twin-Assisted Deep Reinforcement Learning for Interactive Beamforming
To address the aforementioned issues, we propose a DT-assisted offline DRL method in this section, which comprises the DT construction phase and the DT-assisted offline DRL training phase. In the design of our proposed method, both of the two phase leverage pre-collected datasets for training, allowing the agent to learn and improve without the need for ongoing interaction with the environment. The specific designs of these two phase are illustrated in Fig. 4 and analyzed below.
@@ IV-A1 Problem Formulation
where $\mathcal{L}_{D}$ quantifies the distance between the true and approximated distributions. $\hat{p}$ is the approximated distribution. $\theta_{g}$ denotes the parameters in the DT network to be trained.
@@ IV-A3 Network Construction
[…]
where $\theta_{g}$ and $\theta_{d}$ are the parameters of $G$ and $D$, respectively. The process iterates until the Nash equilibrium is achieved, at which point the generator produces data that is indistinguishable from real data to the discriminator. In this equilibrium, the discriminator outputs a probability of 0.5 for all inputs, indicating its inability to distinguish between real and fake data. Detailed hyperparameters are provided in Section V-A.
@@
$\Delta_{e}(\mathbf{s},a)\buildrel\textstyle.\over{=}Q_{e}(\mathbf{s},a)-Q^{*}(\mathbf{s},a),$ (38)
@@
$\|\Delta^{*}(\mathbf{s},a)\|\buildrel\textstyle.\over{=}\lim_{e\to\infty}\|\Delta_{e}(\mathbf{s},a)\|\leq\frac{\xi_{\mathrm{max}}\delta_{\mathcal{T}}}{1-\rho},$ (39)
@@
  • Learning rate $\sum_{e=1}^{\infty}\xi_{e}=\infty$ and $\sum_{e=1}^{\infty}\xi_{e}^{2}<\infty$ uniformly w.p.1., , and $\xi_{\mathrm{max}}=\max_{e}\xi_{e}$.
@@
  • $\delta_{\mathcal{T}}\buildrel\textstyle.\over{=}\sup_{\{\mathbf{s},a\}}\|\mathcal{T}_{\mathrm{DT}}\left(Q^{*}(\mathbf{s},a)\right)-\mathcal{T}_{\mathrm{RW}}\left(Q^{*}(\mathbf{s},a)\right)\|$.
@@
Derived from the training loss, we can derive the error iterative function (40) shown at the bottom of the page. According to the Bellman operator, the second in (40) can be rewritten as (IV-A3). Therefore, according to [42], we have the following inequality when $e\to\infty$:
@@
1 Initialize: Initialize action-value function $Q$ with random weights $\theta_{q}$. Initialize target action-value function $Q^{-}$ with weights $\theta_{q}^{-}=\theta_{q}$. Initialize the cGAN model with random parameters $\theta_{g}$ and $\theta_{d}$.
2 Data Collection: Randomly select actions and interact with the environment with policy $\hat{\pi}_{q}$ to collect some data $(\mathbf{s}_{t},a_{t},r_{t},\mathbf{s}_{t+1})$ and obtain a dataset $\mathcal{D}_{\mathrm{DT}}$.
3 Train the Digital Twin: Train the cGAN model $\theta_{g}$ and $\theta_{d}$ using the loss function (37) and dataset $\mathcal{D}_{\mathrm{DT}}$.
4 for $\bar{e}=1,\ldots,\bar{E}$ do
5       for $\hat{e}=1,\ldots,\hat{E}$ do
6             Update the discriminator parameters $\theta_{d}$ by performing a gradient descent step.
7       end for
8      Update the generator parameters $\theta_{g}$ by performing a gradient descent step with well-trained $\theta_{d}^{*}$.
9 end for
10Train the DRL Agent:
11 for $e=1,\ldots,E$ do
12      Initialize observation state $\mathbf{s}_{1}$.
13       for $t=1,\ldots,T$ do
14            Observe the current state $\mathbf{s}_{t}$.
15             With probability $P_{e}=0.2(1-\frac{e}{E})$ select a random action $a_{t}$. Otherwise select $a_{t}=\text{argmax}_{a}Q(\mathbf{s}_{t},a;\theta)$.
16             Execute beamforming action $a_{t}$ in the digital twin environment, interact with the well-trained generator $G_{\theta_{g}^{*}}$ and obtain the reward $r_{t}$ along with the next state $\mathbf{s}_{t+1}$ according to the output of it.
17             Store transition $(\mathbf{s}_{t},a_{t},r_{t},\mathbf{s}_{t+1})$ in the replay buffer $\Omega$.
18             Randomly sample minibatch of transitions from $\Omega$ and train the Q-network by performing a gradient descent step with loss function (35).
19             Update the Q-network $Q$.
20             Every $C$ steps update the target network $Q^{-}$ with $\theta_{q}^{-}=\theta_{q}$.
21       end for
22
23 end for
Algorithm 1 Digital Twin-Assisted Offline DRL for Beam Selection
@@ V-A Simulation Setup
[…]
The pre-collected dataset used for training DT is obtained through initial agent-environment interactions by 200 episodes. During DT training, the Adam optimizer is used to optimize the gradient of cGAN, with a learning rate of $2\times 10^{-4}$, a batchsize of $64$, and parameters $(\beta_{1},\beta_{2})=(0.5,0.999)$. The number of iterations and epochs for updating the discriminator and generator are set to $\hat{E}=1$ and $\bar{E}=100$, respectively.
The parameters for MDP modeling and DRL training are listed as follows: $\tau_{\mu}=5,b_{0}=5,b_{1}=1,b_{2}=b_{3}=5,\rho=0.99,\alpha_{\mathrm{con}}=0.1,E=1000,T=64,C=200$. The capacity ceiling of the replay buffer is $|\Omega|=5\times 10^{4}$. Adam optimizer is also utilized here with the learning rate and batchsize set to $10^{-4}$ and $16$, respectively. The initial learning rate $\xi$ is set to 0.01 and is then scheduled by the exponential decay with exponential decay factor of 0.9.
@@ V-B Fitting Accuracy of Digital Twin Module
(a) Generator
(b) Discriminator
Figure 5: Convergence behaviors of the cGAN-based DT module.
As shown in Fig. 5, the generator and discriminator converge to their theoretical convergency values of 0.693 and 1.386 [43], respectively, within 100 epochs. This convergence, observed under transmitting power levels ranging from 15 dBmW to 30 dBmW, demonstrates the stability of the DT module. Additionally, Table I presents the values of three metrics across different transmitting power levels. All values remain below 0.02, indicating that the data distribution generated by the digital twin closely approximates the real one. This precise modeling lays a strong foundation for the following DRL assistance. According to the derivation in Theorem 1, the generator’s error directly affects the discrepancy between $\mathcal{T}_{\mathrm{DT}}(Q^{*})$ and $\mathcal{T}_{\mathrm{RW}}(Q^{*})$. Consequently, a smaller generator error leads to a smaller upper bound of Q-error, enabling our proposed method to approach the optimal value.
TABLE I: Distribution differences between DT generated data and real data.
@@ V-E Performance under Different CFAR Constraints
Fig. 10 presents a comparative analysis of tracking steps for different beamforming strategies under varying CFAR constraints, denoted by values of $P_{\mathrm{FA}}$. The four subfigures correspond to $P_{\mathrm{FA}}=10^{-1},10^{-2},10^{-3},$ and $10^{-4}$, illustrating the performance of online DRL, offline DRL, DT-DRL with CQL, and DT-DRL without CQL. It can be observed that as $P_{\mathrm{FA}}$ decreases, tracking steps tend to increase across all methods, indicating a direct impact of CFAR constraints on beamforming performance and a trade-off between detection reliability and tracking efficiency. Stricter false alarm control reduces detection sensitivity, necessitating more steps for beam alignment. Furthermore, DT-DRL with CQL maintains relatively low tracking steps compared to DT-DRL without CQL, confirming that incorporating CQL enhances the efficiency of DT-assisted DRL models. Offline DRL requires more tracking steps than DT-DRL with CQL, underscoring the benefits of the DT module. Finally, DT-DRL with CQL proves to be the most effective approach under all CFAR constraints, reinforcing the advantages of our proposed method for adaptive beamforming.
@@ V-F Performance at Different Velocities of The Moving Target
Taking a step further, Fig. 11 illustrates the impact of dynamic environmental changes on beamforming strategies, with target velocity changing from $v=0$ to $v=20$m/s. The bars represent the number of tracking steps required for beam alignment. It can be seen that as velocity increases, the number of tracking steps required for beam alignment rises in both DT-assisted and non-DT-assisted cases, reflecting the greater challenge of tracking highly mobile targets. However, offline DRL with DT consistently outperforms its counterpart without DT, requiring fewer tracking steps across all velocity levels. This improvement is attributed to DT’s ability to enhance beam tracking performance through augmented data support. The performance gap widens as velocity increases, emphasizing the crucial role of DT in enhancing beam tracking under dynamic conditions. Without DT, the system struggles to adapt efficiently, leading to a sharper increase in tracking steps. These findings suggest that DT-assisted offline DRL excels in fast-changing environments, where rapid adjustments are essential for maintaining effective beam alignment.

### v2

@@ IV Digital Twin-Assisted Deep Reinforcement Learning for Interactive Beamforming
To address the aforementioned issues, we propose a DT-assisted offline DRL method in this section, which comprises the DT construction phase and the DT-assisted offline DRL training phase. In the design of our proposed method, both of the two phases leverage pre-collected datasets for training, allowing the agent to learn and improve without the need for ongoing interaction with the environment. The specific designs of these two phases are illustrated in Fig. 4 and analyzed below.
@@ IV-A1 Problem Formulation
where $\mathcal{L}_{D}$ quantifies the distance between the true and approximated distributions. $\hat{p}$ is the approximated distribution. $\theta_{g}$ denotes the parameters in the DT network to be trained. In this work, $\mathcal{L}_{D}$ is computed by the Jensen-Shannon (JS) divergence.
@@ IV-A3 Network Construction
$\underset{G}{\min}\,\underset{D}{\max}\,\mathbb{E}\!\left[\log D_{\theta_{d}}\!\left(P_{D}^{t+1}\right)\!+\!\log\left(1\!-\!D_{\theta_{d}}\!\left(G_{\theta_{g}}\!\left(\mathbf{z}|\mathbf{u}_{t}\right)\right)\right)\right],$
where $\theta_{g}$ and $\theta_{d}$ are the parameters of $G$ and $D$, respectively. The process iterates until the Nash equilibrium is achieved, at which point the generator produces data that is indistinguishable from real data to the discriminator. In this equilibrium, the discriminator outputs a probability of 0.5 for all inputs, indicating its inability to distinguish between real and fake data. According to [13, 40], such a balance is equivalent to minimizing the JS divergence between the approximated distribution $\hat{p}(P_{D}^{t+1}|\mathbf{u}_{t})$ and true distribution $p(P_{D}^{t+1}|\mathbf{u}_{t})$. Detailed hyperparameters are provided in Section V-A.
@@
$\Delta_{e}(\mathbf{s},a)\mathrel{\mathop{\kern 0.0pt=}\limits^{\textstyle.}}Q_{e}(\mathbf{s},a)-Q^{*}(\mathbf{s},a),$ (40)
@@
$\|\Delta^{*}(\mathbf{s},a)\|\mathrel{\mathop{\kern 0.0pt=}\limits^{\textstyle.}}\lim_{e\to\infty}\|\Delta_{e}(\mathbf{s},a)\|\leq\frac{\xi_{\mathrm{max}}\delta_{\mathcal{T}}}{1-\rho},$ (41)
@@
  • Learning rate $\sum_{e=1}^{\infty}\xi_{e}=\infty$ and $\sum_{e=1}^{\infty}\xi_{e}^{2}<\infty$ uniformly w.p. 1, and $\xi_{\mathrm{max}}=\max_{e}\xi_{e}$.
@@
  • $\delta_{\mathcal{T}}\mathrel{\mathop{\kern 0.0pt=}\limits^{\textstyle.}}\sup_{\{\mathbf{s},a\}}\|\mathcal{T}_{\mathrm{DT}}\left(Q^{*}(\mathbf{s},a)\right)-\mathcal{T}_{\mathrm{RW}}\left(Q^{*}(\mathbf{s},a)\right)\|$.
@@
Derived from the training loss, we can derive the error iterative function (42) shown at the top of the page. According to the Bellman operator, the second term in (42) can be rewritten as (IV-A3). Therefore, according to [28], we have the following inequality when $e\to\infty$:
@@
From a physical perspective, Theorem 1 implies that the proposed learning scheme achieves stable beam selection performance even under modeling discrepancies between the DT and the real environment, as the Q-value deviation remains bounded within a finite margin during convergence.
@@ IV-B The DT-assisted offline DRL training
Once the offline training is completed, the DT is no longer required during practical implementation. The trained policy, as formulated in (37), can then be directly deployed.
Initialize: Initialize action-value function $Q$ with random weights $\theta_{q}$. Initialize target action-value function $Q^{-}$ with weights $\theta_{q}^{-}=\theta_{q}$. Initialize the cGAN model with random parameters $\theta_{g}$ and $\theta_{d}$.
Data Collection: Randomly select actions and interact with the environment with policy $\hat{\pi}_{q}$ to collect some data $(\mathbf{s}_{t},a_{t},r_{t},\mathbf{s}_{t+1})$ and obtain a dataset $\mathcal{D}_{\mathrm{DT}}$.
Train the Digital Twin: Train the cGAN model $\theta_{g}$ and $\theta_{d}$ using the loss function (39) and dataset $\mathcal{D}_{\mathrm{DT}}$.
for $\bar{e}=1,\ldots,\bar{E}$ do
 for $\hat{e}=1,\ldots,\hat{E}$ do
      Update the discriminator parameters $\theta_{d}$ via gradient ascent on: $\max_{\theta_{d}}\mathbb{E}[\log D_{\theta_{d}}(P_{D}^{t+1})+\log(1-D_{\theta_{d}}(G_{\theta_{g}}(\mathbf{z}|\mathbf{u}_{t})))].$
   end for
  Update the generator parameters $\theta_{g}$ via gradient descent with well-trained $\theta_{d}^{*}$: $\min_{\theta_{g}}\mathbb{E}[\log(1-D_{\theta_{d}^{*}}(G_{\theta_{g}}(\mathbf{z}|\mathbf{u}_{t})))].$
end for
Train the DRL Agent:
for $e=1,\ldots,E$ do
  Initialize observation state $\mathbf{s}_{1}$.
 for $t=1,\ldots,T$ do
     Observe the current state $\mathbf{s}_{t}$.
      With probability $P_{e}=0.2(1-\frac{e}{E})$ select a random action $a_{t}$. Otherwise select $a_{t}=\text{argmax}_{a}Q(\mathbf{s}_{t},a;\theta)$.
      Execute beamforming action $a_{t}$ in the digital twin environment, interact with the well-trained generator $G_{\theta_{g}^{*}}$ and obtain the reward $r_{t}$ along with the next state $\mathbf{s}_{t+1}$ according to the output of it.
      Store transition $(\mathbf{s}_{t},a_{t},r_{t},\mathbf{s}_{t+1})$ in the replay buffer $\Omega$.
      Randomly sample minibatch of transitions from $\Omega$ and train the Q-network by performing a gradient descent step with loss function (36).
      Update the Q-network $Q$.
      Every $C$ steps update the target network $Q^{-}$ with $\theta_{q}^{-}=\theta_{q}$.
   end for
end for
Algorithm 1 Digital Twin-Assisted Offline DRL for Beam Selection
IV-C Computational Complexity Analysis
The computational complexity of the cGAN-based DT mainly arises from the fully connected (FC) layers in both the generator and the discriminator. Let $C_{in}^{(l)}$ and $C_{out}^{(l)}$ denote the input and output dimensions of the $l$-th FC layer, respectively, and let $L_{G}$ and $L_{D}$ represent the total number of layers in the generator and discriminator. The computational complexity of a forward pass can thus be expressed as $\mathcal{O}\left(\sum_{l=1}^{L_{G}}C_{in,G}^{(l)}C_{out,G}^{(l)}+\sum_{l=1}^{L_{D}}C_{in,D}^{(l)}C_{out,D}^{(l)}\right)$. According to the network architecture shown in Fig. 3, we have $L_{G}=3,\,C_{in,G}^{(1)}=d_{in,G},\,C_{out,G}^{(1)}=C_{in,G}^{(2)}=128,\,C_{out,G}^{(2)}=C_{in,G}^{(3)}=64,\,C_{out,G}^{(3)}=1;\,L_{D}=3,\,C_{in,D}^{(1)}=d_{in,D},\,C_{out,D}^{(1)}=C_{in,D}^{(2)}=64,\,C_{out,D}^{(2)}=C_{in,D}^{(3)}=128,\,C_{out,D}^{(3)}=1$, where $d_{in,G}$ and $d_{in,D}$ denote the input dimensions of generator and discriminator, respectively. Since the proposed network employs lightweight FC layers with moderate hidden dimensions, the overall computational cost grows linearly with the input dimension. Moreover, the computational overhead of the DT only exists in the offline training phase. Once the twin environment is constructed, no additional computational burden is introduced during actual engineering implementation.
The computational complexity of the DRL module is evaluated in a similar way. During both training episodes and online inference, the dominant computational cost arises from matrix multiplications within the neural network layers. Let $L_{Q}$ represent the total number of layers in the dueling Q-network, the computational complexity can be expressed as $\mathcal{O}\left(\sum_{l=1}^{L_{Q}}C_{in,Q}^{(l)}C_{out,Q}^{(l)}\right)$. According to the network structure illustrated in Fig. 3, $L_{Q}=3,\,C_{in,Q}^{(1)}=d_{in,Q},\,C_{out,Q}^{(1)}=C_{in,Q}^{(2)}=32,\,C_{out,Q}^{(2)}=C_{in,Q}^{(3)}=16,\,C_{out,Q}^{(3)}=2\tau_{\mu},$ where $d_{in,Q}$ represents the input dimension.
@@ V-A Simulation Setup
TABLE I: Parameters for DT and DRL training.
Task
Batchsize Epoch / Iteration Optimizer Learning Rate Network Depth Others
DT Training
64 $\bar{E}=100,\hat{E}=1$ Adam (0.5, 0.999) $2\times 10^{-4}$ $L_{G},L_{D}=3$ -
DRL Training
16 $E=1000,T=64,C=200$ Adam (0.9, 0.999) $10^{-4}\cdot 0.9^{e}$ $L_{Q}=3$ $\alpha_{\mathrm{con}}=0.1,|\Omega|=5\times 10^{4}$
We consider a cell-free ISAC scenario with a transmit AP and $N=4$ receive APs distributed in the area of size 400m $\times$ 400m. To be exact, they are positioned at coordinates (250m, 250m), (100m, 100m), (100m, 400m), (400m, 100m), and (400m, 400m), respectively, ensuring equidistant placement. Each AP is equipped with $M=64$ antennas. The target trajectory follows a piecewise-linear motion model, with the target starting from a random initial position uniformly distributed within the APs’ coverage area and moves randomly at a velocity of $v=0\sim 5\text{m/s}$. The beamforming codebook uniformly quantizes the azimuth range of $0,2\pi)$ into 64 discrete beams. The transmitting power, the CFAR constraint $P_{\mathrm{FA}}$, and the detection probability threshold $\zeta$ are set to 20 dBmW, $10^{-3}$, and 0.9, respectively, unless explicitly stated otherwise.
The large-scale fading coefficient $\beta_{ij}$ between objects $i$ and $j$ is modeled to capture the effects of path loss and is formulated as: $\beta_{ij}\mathrm{[dB]}=-34.5-\alpha\cdot 10\log{10}\left(\frac{d_{ij}}{d_{0}}\right)$, where $d_{ij}$ denotes the Euclidean distance between objects $i$ and $j$, $d_{0}=1\text{m}$ is the reference distance, and the path loss exponent is set to $\alpha=3.7$. Moreover, the power of the AWGN is computed as $\sigma^{2}\mathrm{[dB]}=-174+n_{f}+10\log_{10}\left(B\right)$, where $n_{f}=9\text{dB}$ is the noise figure, and $B=20\text{MHz}$ represents the signal bandwidth.
The pre-collected dataset used for training DT is obtained through initial agent-environment interactions by 200 episodes. The parameters for MDP modeling are listed as follows: $\tau_{\mu}=5,b_{0}=5,b_{1}=1,b_{2}=b_{3}=5,\rho=0.99$. Other parameters for DT and DRL training are summarized in Tabel I, where $|\Omega|$ denotes the capacity ceiling of the replay buffer.
@@ V-B Fitting Accuracy of Digital Twin Module
As shown in Fig. 5, the generator and discriminator converge to their theoretical convergency values of 0.693 and 1.386 [13], respectively, within 100 epochs. This convergence, observed under transmitting power levels ranging from 15 dBmW to 30 dBmW, demonstrates the stability of the DT module. The abrupt loss fluctuation observed around epoch 50, particularly for TP = 30 dBmW, arises from the intrinsic adversarial dynamics of conditional GAN training. In such a minimax setting, temporary imbalances between the discriminator and generator can lead to sudden spikes or drops in the loss curves. This phenomenon is more evident under high-TP conditions, where stronger signals make the data distribution initially easier for the discriminator to classify, causing a temporary dominance that increases the generator loss. As training progresses, the generator adapts to restore equilibrium, producing the observed oscillatory behavior.
Additionally, Table II presents the values of three metrics across different transmitting power levels. All values remain below 0.02, indicating that the data distribution generated by the digital twin closely approximates the real one. This precise modeling lays a strong foundation for the following DRL assistance. According to the derivation in Theorem 1, the generator’s error directly affects the discrepancy between $\mathcal{T}_{\mathrm{DT}}(Q^{*})$ and $\mathcal{T}_{\mathrm{RW}}(Q^{*})$. Consequently, a smaller generator error leads to a smaller upper bound of Q-error, enabling our proposed method to approach the optimal value.
TABLE II: Distribution differences between DT generated data and real data.
@@
(a) Generator
(b) Discriminator
Figure 5: Convergence behaviors of the cGAN-based DT module.
@@ V-C Performance Comparisons with Other Methods
Furthermore, we perform additional numerical simulations to investigate the impact of the parameter $b_{2}$ on the agent’s learning behavior. We observe that a moderate value of $b_{2}$ enhances the information-capturing ability of the shaping function, enabling the agent to extract meaningful feedback in the early stage while still pursuing improved performance in later stages. In contrast, an excessively small $b_{2}$ weaken the reward shaping effect, preventing the agent from learning effectively. Conversely, an excessively large $b_{2}$ causes the reward to saturate too quickly, leading the agent to prematurely settle for suboptimal detection probabilities rather than continuing to explore better strategies. As a result, both extremes of $b_{2}$ severely degrade the overall convergence and final performance, whereas moderate values strike the desired balance between exploration and stability.
@@ V-E Performance under Different CFAR Constraints
Fig. 10 presents a comparative analysis of tracking steps for different beamforming strategies under varying CFAR constraints, denoted by values of $P_{\mathrm{FA}}$. The four subfigures correspond to $P_{\mathrm{FA}}=10^{-1},10^{-2},10^{-3},$ and $10^{-4}$, illustrating the performance of online DRL, offline DRL, DT-DRL with CQL, and DT-DRL without CQL, respectively. It can be observed that, as $P_{\mathrm{FA}}$ decreases, tracking steps tend to increase across all methods, indicating a direct impact of CFAR constraints on beamforming performance and a trade-off between detection reliability and tracking efficiency. Stricter false alarm control reduces detection sensitivity, necessitating more steps for beam alignment. Furthermore, DT-DRL with CQL maintains relatively fewer tracking steps compared to DT-DRL without CQL, confirming that incorporating CQL enhances the efficiency of DT-assisted DRL models. Offline DRL requires more tracking steps than DT-DRL with CQL, underscoring the benefits of the DT module. Finally, DT-DRL with CQL proves to be the most effective approach under all CFAR constraints, reinforcing the advantages of our proposed framework for adaptive beamforming.
@@ V-F Performance at Different Velocities of The Moving Target
Taking a step further, Fig. 11 illustrates the impact of dynamic environmental changes on beamforming strategies, with target velocity changing from $v=0$ to v=20v=20m/s. The bars represent the number of tracking steps required for beam alignment. It can be seen that as velocity increases, the number of tracking steps required for beam alignment rises in both DT-assisted and non-DT-assisted cases, reflecting the greater challenge of tracking highly mobile targets. However, offline DRL with DT consistently outperforms its counterpart without DT, requiring fewer tracking steps across all velocity levels. This improvement is attributed to the DT’s ability to enhance beam tracking performance through augmented data support. By enriching the diversity of state-action pairs, the DT effectively expands the training domain beyond the original data. Such a capability allows the agent to overcome the limitations of historical data and to generate plausible transition samples corresponding to previously unseen target trajectories or higher target velocities. The performance gap widens as velocity increases, emphasizing the crucial role of DT under dynamic conditions. Without DT, the system struggles to adapt efficiently, leading to a sharper increase in tracking steps. These findings suggest that DT-assisted offline DRL excels in fast-changing environments, where rapid adjustments are essential for maintaining effective beam alignment.